OpenStudy (anonymous):

help please Solve the following system of equations. x + y = 3 2x + 2y = 6

OpenStudy (igreen):

We can solve it using substitution: x + y = 3 2x + 2y = 6 First re-arrange the 2nd equation: x + y = 3 Subtract y to both sides: x = -y + 3 Now we can plug in -y + 3 into the 2nd equation for 'x': 2x + 2y = 6 2(-y + 3) + 2y = 6 Distribute 2 into the parenthesis: -2y + 6 + 2y = 6 Can you add -2y + 2y? @brianna1000

OpenStudy (anonymous):

would it be 0?

OpenStudy (igreen):

Yes, which gives us: 6 = 6 So the answer here will be 'Infinitely many solutions', meaning they're the same line.

OpenStudy (igreen):

Welcome to OpenStudy! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

thank you an i did :)

OpenStudy (igreen):

No problem.